Roofers need to move equipment onto the roof of a building. They have a machine that will pull a platform at varying accelerations. The rope used to pull the platform will snap if a force of 5550 Newtons is exceeded. The platform itself has a mass of 50.0kg. The equipment has a mass of 475 kg. What is the maximum acceleration the machine could lift the equipment without snapping the rope?



1
Expert's answer
2020-12-25T14:07:48-0500
"(m+M)(a+g)=F\\(50+475)(a+9.8)=5550\\a=0.771\frac{m}{s^2}"